# ODMixer: Fine-grained Spatial-temporal MLP for Metro Origin-Destination Prediction

### v2

@@ Abstract
Metro Origin-Destination (OD) prediction is a crucial yet challenging spatial-temporal prediction task in urban computing, which aims to accurately forecast cross-station ridership for optimizing metro scheduling and enhancing overall transport efficiency. Analyzing fine-grained and comprehensive relations among stations effectively is imperative for metro OD prediction. However, existing metro OD models either mix information from multiple OD pairs from the station’s perspective or exclusively focus on a subset of OD pairs. These approaches may overlook fine-grained relations among OD pairs, leading to difficulties in predicting potential anomalous conditions. To address these challenges, we analyze traffic variations from the perspective of all OD pairs and propose a fine-grained spatial-temporal MLP architecture for metro OD prediction, namely ODMixer. Specifically, our ODMixer has double-branch structure and involves the Channel Mixer, the Multi-view Mixer, and the Bidirectional Trend Learner. The Channel Mixer aims to capture short-term temporal relations among OD pairs, the Multi-view Mixer concentrates on capturing relations from both origin and destination perspectives. To model long-term temporal relations, we introduce the Bidirectional Trend Learner. Extensive experiments on two large-scale metro OD prediction datasets HZMOD and SHMO demonstrate the advantages of our ODMixer. Our code is available at https://github.com/KLatitude/ODMixer.
@@ I Introduction
Compared to general traffic OD prediction, metro OD prediction has unique challenges. First, due to varying time intervals between passengers entering and exiting stations, it is impossible to count the orders of passengers who have not yet exited when calculating the OD matrix at any given time, resulting in an incomplete OD matrix. This incompleteness is a significant characteristic of metro OD prediction. Secondly, the large number of metro stations leads to a high-dimensional OD matrix, posing substantial challenges to model design. Additionally, the uneven distribution of metro traffic causes significant differences between OD pairs, resulting in a sparse OD matrix. Analyzing these traffic differences is challenging.
@@
(b) Interval difference between OD pairs
@@
Figure 2: Comparison of flow changes between OD pairs with the same origin. The x-axis represents July 4 and 5, 2019, the y-axis represents the flow value between OD pairs. The OD:(190,196) in the legend means the origin is station 190 and the destination is station 196. Figure (a) shows that although the OD pairs with the same origin have similar change trends, the flow values are very different. Figure (b) illustrates that the flow rate trends of two OD pairs with the origin exhibit significant differences over time.
Over time, various deep learning models have been employed to address the metro OD prediction problem [5, 6, 7, 8]. However, as illustrated in Fig. 1(a), these models predominantly adopt a station view, encoding all data from each station at every time moment in a coarse-grained manner. This approach results in the mixing of data from various OD pairs, causing the encoded features to encompass information from different OD pairs. Consequently, these models face challenges in discerning the flow changes of OD pairs. As shown in Figure 2(a), although the time variation trends of the flow rates of two OD pairs with the origin are similar, the difference in flow rates is very pronounced. Additionally, as shown in Figure 2(b), the time variation trends between the two OD pairs are completely different. Therefore, if the OD matrix is modeled based on the perspective of the stations, the differences between different OD pairs may be ignored, making the model unable to effectively capture the temporal changes in flow, resulting in a decrease in model performance. In reality, given that the purpose of most people’s trips is clear, and the origin and destination of the majority of trips are predetermined, it is more suitable to encode the OD matrix from the perspective of OD pairs. The OD pair view encodes each OD pair independently to distinguish different OD pairs, which is more fine-grained than the station view that mixes multiple OD pairs.
@@
Actually, there exists some models that predict the metro OD matrix from the OD pair view [9, 10], but they only focus a subset of OD pairs in the metro. This constrained approach neglects the richness of information among all OD pairs, resulting in the model’s inability to learn fine-grained relations across the entirety of OD pairs. Consequently, it hampers the model’s capability to address potential emergencies, comprehend the dynamics of traffic changes, making it difficult to provide effective support for the management of the metro. To achieve more comprehensive metro OD prediction, it is imperative to consider information from all OD pairs. However, considering the number of OD pairs as $N^{2}$, the time complexity needed to calculate the relations among $N^{2}$ OD pairs using Transformer or to construct a virtual graph among OD pairs and subsequently employ Graph Convolution Network (GCN) for information aggregation is $O(N^{4})$, as shown in Fig. 1(b). Consequently, keeping the balance between model’s parameters and computation time while considering all OD pairs remains challenging.
Inspired by the efficiency of the MLP architecture [11, 12], we propose a simple yet efficient fine-grained spatial-temporal MLP model named ODMixer, to tackle the aforementioned challenges. Initially, we individually encode the time series of OD pairs base on OD pair view. Subsequently, to capture the intricate relations among OD pairs, we introduce two modules, Channel Mixer and Multi-view Mixer. The Channel Mixer module facilitates the modeling of temporal correlation by calculating relations among features of OD pairs in the temporal dimension. Considering the substantial time required for direct computation of relations among all OD pairs, as with methods like Transformer, and the potential confusion and difficulty in distinguishing important associations, the Multi-view Mixer calculates relations among different OD pairs from two perspectives: Origin and Destination. This enhances the comprehensive understanding of OD pair relations, thereby improving the efficiency of modeling fine-grained relations among OD pairs. Additionally, to exploit the temporal periodicity of metro traffic, we introduce the Bidirectional Trend Learner that perceives the long-term temporal relations between traffic flow and enhance the temporal attributes of features. Our contributions are summarized as follows:
@@ II-A Traffic Flow Prediction
With the rapid development of deep learning, neural networks have been used to solve traffic prediction problems. These representative works include many structural types, mainly including CNN [16, 17], RNN [18, 19], Attention [20, 21], etc. DeepSD[22] proposed an end-to-end deep learning framework and integrates external environment information to realize the automatic discovery of complex supply-demand patterns from the car-hailing service data. Based on three temporal relations, MDL[23] used two three-stream fully convolutional networks (3S-FCNs) to achieve flow prediction for points and edges. DeepCrowd[24] divided the area into multiple fine-grained grids, and combines ConvLSTM and attention mechanisms to analyze and fuse flow information at multiple times to achieve Large-Scale Citywide Crowd Density and Flow Prediction.
Later, the GCN was popular to process the relation between nodes because of the natural network structure in the traffic flow prediction problem [25, 26, 27]. STDGRL[28] designed a spatio-temporal dynamic graph relation learning model. The spatial relation is captured dynamically by dynamically learning the graph structure between nodes, and the long-term temporal relation is learned by using attention. STWave+[29] utilized decoupling technology to partition one-hour traffic data into stable trends and fluctuating events. This approach employs a dual-channel spatiotemporal network to separately model these components, leveraging self-supervised learning and contrastive loss to propagate long-term trend information to hourly trends. Ultimately, it forecasts forthcoming traffic conditions through adaptive fusion techniques.
@@ II-B Origin-Destination Prediction
OD prediction [30] involves the accurate estimation of traffic flow between two regions over a given period of time, presenting a formidable challenge due to its intricate and dynamic nature. As a pivotal aspect of urban traffic management, OD prediction has garnered extensive attention.
As a special type of OD prediction, the metro OD prediction [6, 31, 5, 32, 7, 9, 33] presents unique challenges compared to the general OD prediction, primarily because the precise destinations of passengers are not accurately known until they reach their destinations. This leads to incomplete data at the current moment, making it challenging to effectively depict the real-time traffic distribution. To address this issue and obtain more comprehensive data on traffic distribution, several studies focused on predicting unfinished orders to supplement the incomplete OD matrix. For instance, in HIAM[5], long and short historical distributions were utilized to estimate the distribution of unfinished orders at the current time. This study integrated the completed and unfinished distributions in the feature space. In Gong et al.[6], an indication matrix was created, with values assigned based on the assumption that the travel time of each OD pair at each timestamp follows a normal distribution. However, this approach overlooked the unfinished OD pair data.
@@ II-C MLP-based Model
Figure 3: (a) illustrates the architecture of double-branch ODMixer, both branches learn features using ODIM. Subsequently, the two branches interact with each other using the BTL, contributing to the final output. (b) depicts the BTL module, which aims to jointly model the features from both branches, thereby enhancing the information exchange. (c) represents the ODIM module, which models the temporal attributes of the OD pair features using the Channel Mixer, while the comprehensive relations among OD pairs are learned using the Multi-view Mixer. (d) shows the Channel Mixer, Origin Mixer, Des Mixer.
@@ III Preliminaries
Definition 1. The variable $\mathrm{IOD}_{t}^{d}(i,j)$ is defined as the quantity of traffic entering from station $i$ and exiting from station $j$ at the $t^{th}$ time interval of the $d^{th}$ day.
Definition 2. The notation $\mathrm{UOD}_{t}^{d}(i,j)$ signifies the flow entering station $i$ at the $t^{th}$ time interval and exiting from station $j$ after the $t^{th}$ time interval of the $d^{th}$ day.
Definition 3. The notation $\mathrm{OD}_{t}^{d}(i,j)$ denotes the flow at time $t$ entering station $i$ and subsequently exiting from station $j$ at and after the $t^{th}$ time interval of the $d^{th}$ day. $\mathrm{OD}_{t}^{d}=\mathrm{IOD}_{t}^{d}+\mathrm{UOD}_{t}^{d}$.
Definition 4. The notation $\mathrm{unf}_{t}^{d}(i)$ represents the flow at the $t^{th}$ time interval entering station $i$ but not exiting from any station at the $t^{th}$ time interval of the $d^{th}$ day. $\mathrm{unf}_{t}^{d}(i)=\sum_{j=1}^{N}\mathrm{UOD}_{t}^{d}(i,j)$.
@@ IV Methodology
$\mathrm{unf}_{t}^{d}\in\mathbb{R}^{N}$ the unfinished order vector
$\mathrm{OD}_{t}^{d}\in\mathbb{R}^{N\times N}$ the complete OD matrix
$\mathrm{UOD}_{t}^{d}\in\mathbb{R}^{N\times N}$ the OD matrix for unfinished orders
$\mathrm{IOD}_{t}^{d}\in\mathbb{R}^{N\times N}$ the OD matrix for finished orders
@@ IV-A Architecture Overview
IV-B OD Matrix Processing
Since ODMixer contains two branches (prev and cur), the input of the prev branch is the OD matrix at the same time yesterday. Since yesterday’s order information has been completed, its input is a complete OD matrix and no operation is required. The input of the cur branch is the OD matrix at the current time, so it is incomplete and needs to be processed. In this subsection, we introduce the approach for estimating the current UOD distribution by leveraging the historical UOD distribution and the fine-grained encoding of the OD matrix from the OD pair perspective, using the Embedding Layer.
IV-B1 Unfinished Orders Processing
@@
Acknowledging the temporal similarity characteristic of metro traffic, where the current OD traffic distribution closely resembles that of the previous time, we estimate the current UOD distribution by referencing historical UOD distributions. Our approach specifically incorporates two temporal scales: long-term and short-term UOD distributions. The calculations are outlined as follows:
$\mathrm{\widetilde{UOD}}=\frac{\mathrm{unf}\ast(\mathrm{UOD}_{S}+\mathrm{UOD}_{L})}{2}$ (2)
Here, $\mathrm{\widetilde{UOD}}\in\mathbb{R}^{T\times N\times N}$ represents an estimation of unfinished orders over T time intervals, $\mathrm{unf}\in\mathbb{R}^{T\times N}$ is the number of unfinished orders of all stations, the symbol $\ast$ denotes the multiplication of the corresponding elements. $\mathrm{UOD}_{S}\in\mathbb{R}^{T\times N\times N}$ represents the short-term UOD distribution, corresponding to the UOD distribution at the same time yesterday, and $\mathrm{UOD}_{L}\in\mathbb{R}^{T\times N\times N}$ represents the long-term UOD distribution, aligned with the UOD distribution observed at the corresponding time on the same day of the previous week. This computation aims to generate a more comprehensive and accurate estimate of the UOD distribution at the current time by combining both short-term and long-term UOD distributions.
@@
$\mathrm{OD}=\mathrm{IOD}+\mathrm{\widetilde{UOD}}$ (3)
where $\mathrm{OD}\in\mathbb{R}^{T\times N\times N}$ is the complete OD matrix at the current time, $\mathrm{IOD}\in\mathbb{R}^{T\times N\times N}$ represents the matrix of actual completed orders, and $\mathrm{\widetilde{UOD}}$ denotes the matrix of unfinished orders computed from the historical short-term and long-term UOD distributions. This operation ensures that the model receives more comprehensive and accurate input information, improving passenger destination prediction by combining information from both completed and unfinished orders.
IV-B2 Embedding Layer

### v4

@@ Abstract
Metro Origin-Destination (OD) prediction is a crucial yet challenging spatial-temporal prediction task in urban computing, which aims to accurately forecast cross-station ridership for optimizing metro scheduling and enhancing overall transport efficiency. Analyzing fine-grained and comprehensive relations among stations effectively is imperative for metro OD prediction. However, existing metro OD models either mix information from multiple OD pairs from the station’s perspective or exclusively focus on a subset of OD pairs. These approaches may overlook fine-grained relations among OD pairs, leading to difficulties in predicting potential anomalous conditions. To address these challenges, we learn traffic evolution from the perspective of all OD pairs and propose a fine-grained spatial-temporal MLP architecture for metro OD prediction, namely ODMixer. Specifically, our ODMixer has double-branch structure and involves the Channel Mixer, the Multi-view Mixer, and the Bidirectional Trend Learner. The Channel Mixer aims to capture short-term temporal relations among OD pairs, the Multi-view Mixer concentrates on capturing spatial relations from both origin and destination perspectives. To model long-term temporal relations, we introduce the Bidirectional Trend Learner. Extensive experiments on two large-scale metro OD prediction datasets HZMOD and SHMO demonstrate the advantages of our ODMixer. Our code is available at https://github.com/KLatitude/ODMixer.
@@ I Introduction
Metro OD prediction poses unique challenges compared to general OD prediction [5, 6, 7]. First, the time gap between passenger entry and exit often causes these events to occur in different time intervals. As a result, the OD matrix for the current time interval cannot fully represent passengers still within the system, leading to incompleteness. Second, the high number of metro stations results in a high-dimensional OD matrix, complicating model design. Additionally, uneven passenger flow distribution creates significant disparities between OD pairs, contributing to matrix sparsity. These factors make metro OD matrix contain highly complex spatiotemporal relations, making metro OD prediction exceptionally challenging.
For metro OD prediction, recent years have seen the rise of deep learning models leveraging powerful nonlinear modeling capabilities [6, 7, 8, 9]. These models employ architectures such as RNNs, TCNs, and Transformers to capture temporal relations, while GCNs and Transformers are commonly utilized to learn spatial dependencies. However, most of these approaches adopt a station-based perspective, as illustrated in Fig. 1(a). This method encodes data for each station during each time interval, causing the flow data of different OD pairs with the same origin to be mixed at the input layer. As a result, the extracted features combine information from multiple OD pairs, making it challenging for the model to distinguish between them when learning spatiotemporal relations. This limitation makes the station-based perspective a coarse-grained modeling method. As depicted in Fig. 2(a), while the overall flow trends of two OD pairs with the same origin may appear similar over time, significant differences in their flow values, particularly during peak hours, are evident. Moreover, in some cases, as shown in Fig. 2(b), the flow trends of OD pairs with the same origin can be entirely opposite. Mixing these OD pairs at the input layer overlooks such differences, hindering the model’s ability to effectively learn spatiotemporal relations. To address this, we propose an OD pair-based modeling perspective, as shown in the Fig. 1(a). This approach models each OD pair independently, allowing the model to differentiate between OD pairs and accurately capture their complex spatiotemporal relations. Compared to the station-based perspective, the OD pair-based modeling approach provides a fine-grained representation of the OD matrix, enabling the model to learn more fine-grained and accurate patterns in the data. Additionally, from Fig. 2(a) and (b), we can see that for each OD pair, the OD traffic at adjacent day exhibits a very similar trend. However, existing methods [6] do not consider this similarity and fail to incorporate historical traffic information, leading to sub-optimal model performance.
@@
(b) Trend difference between OD pairs
@@
Figure 2: Comparison of flow changes between OD pairs with the same origin. The x-axis represents July 4 and 5, 2019, the y-axis represents the flow value between OD pairs. The OD: (190,196) in the legend means the origin is station 190 and the destination is station 196. Figure (a) shows that although the OD pairs with the same origin have similar change trends, the flow values are very different. Figure (b) illustrates that the flow rate trends of two OD pairs with the same origin exhibit significant differences over time.
@@
Actually, there exist some models that predict the metro OD matrix from the OD pair view [10, 11], but they only focus on a subset of OD pairs in the metro. This constrained approach neglects the richness of information among all OD pairs, resulting in the model’s inability to learn fine-grained relations across the entirety of OD pairs. Consequently, it hampers the model’s capability to address potential emergencies and comprehend the dynamics of traffic changes, making it difficult to provide effective support for metro management. To achieve more comprehensive metro OD prediction, it is imperative to consider information from all OD pairs. However, considering the number of OD pairs as $N^{2}$, the time complexity needed to calculate the relations among $N^{2}$ OD pairs using Transformer or to construct a virtual graph among OD pairs and subsequently employ Graph Convolution Network (GCN) for information aggregation is $O(N^{4})$, as shown in Fig. 1(b). Therefore, balancing the model’s parameters and computation time while considering all OD pairs remains challenging.
Inspired by the efficiency of the MLP architecture [12, 13], we propose ODMixer, a simple yet effective fine-grained spatiotemporal MLP architecture. Unlike existing metro OD prediction models that adopt coarse-grained station-based modeling, ODMixer employs a fine-grained OD pair-based modeling approach to comprehensively capture the complex spatial-temporal relations in metro OD prediction. To address the issue of incomplete input data in metro OD prediction, we introduce the OD Matrix Process module, which leverages periodicity to fill missing values in the OD matrix. Once the OD matrix is complete, we independently model each OD pair from an OD pair-based perspective. This approach not only considers each OD pair separately but also mitigates the challenges posed by uneven distribution in the OD matrix. To capture the intricate spatiotemporal relationships within the OD matrix, we design specialized modules. For temporal relations, we account for both short-term and long-term dependencies. The Channel Mixer operates on the most recent features of each OD pair to capture short-term temporal relations. To model long-term trends, such as periodicity inherent in metro traffic, we employ a dual-branch structure coupled with a Bidirectional Trend Learner. Regarding spatial relations, we treat each OD pair as a node. Given the computational challenges posed by models like Transformers and GCNs in processing all OD pairs, we adopt a decomposition strategy. Specifically, we decompose spatial relations based on the origin and destination perspectives and design the Multi-view Mixer module. This module separately calculates spatial relations from the origin and destination perspectives, utilizing a Mixer structure for each. This design enhances the model’s efficiency and its ability to learn spatial relationships across all OD pairs. Our contributions are summarized as follows:
@@ II-A Traffic Flow Prediction
With the rapid development of deep learning, neural networks have been used to solve traffic prediction problems. These representative works include many structural types, mainly including CNN [17, 18], RNN [19, 20], Attention [21, 22], etc. DeepSD[23] proposed an end-to-end deep learning framework and integrates external environment information to realize the automatic discovery of complex supply-demand patterns from the car-hailing service data. Based on three temporal relations, MDL[24] used two three-stream fully convolutional networks (3S-FCNs) to achieve flow prediction for points and edges. DeepCrowd[25] divided the area into multiple fine-grained grids, and combines ConvLSTM and attention mechanisms to analyze and fuse flow information at multiple times to achieve Large-Scale Citywide Crowd Density and Flow Prediction. Later, the GCN was popular to process the relation between nodes because of the natural network structure in the traffic flow prediction problem [26, 27, 28]. STDGRL[29] designed a spatio-temporal dynamic graph relation learning model. STWave+[30] utilized decoupling technology to partition one-hour traffic data into stable trends and fluctuating events. This approach employs a dual-channel spatiotemporal network to separately model these components, leveraging self-supervised learning and contrastive loss to propagate long-term trend information to hourly trends.
@@ II-B Origin-Destination Prediction
OD prediction [31, 32, 33, 34] involves the accurate estimation of traffic flow between two regions over a given period of time, presenting a formidable challenge due to its intricate and dynamic nature. As a pivotal aspect of urban traffic management, OD prediction has garnered extensive attention.
As a special type of OD prediction, the metro OD prediction [7, 35, 6, 36, 8, 10, 37, 5] presents unique challenges compared to the general OD prediction, primarily because the precise destinations of passengers are not accurately known until they reach their destinations. This leads to incomplete data at the current moment, making it challenging to effectively depict the real-time traffic distribution. To address this issue and obtain more comprehensive data on traffic distribution, several studies focused on predicting unfinished orders to supplement the incomplete OD matrix. For instance, in HIAM[6], long and short historical distributions were utilized to estimate the distribution of unfinished orders at the current time. This study integrated the completed and unfinished distributions in the feature space. Gong et al. [7] created an indication matrix, with values assigned based on the assumption that the travel time of each OD pair at each timestamp follows a normal distribution. However, this approach overlooked the unfinished OD pair data. MVPF [5] proposed a multi-view passenger flow evolution trend based OD matrix prediction method that combines individual station and cross-station learning, uses GRU and EGAT models to learn the spatial-temporal-dependent representation of stations, and defines a transfer matrix to capture passenger mobility patterns.
@@ II-C MLP-based Model
Figure 3: The overall framework of the ODMixer and its essential modules. (a) illustrates the architecture of double-branch ODMixer, both branches learn features using ODIM. Subsequently, the two branches interact with each other using the BTL, contributing to the final output. (b) depicts the BTL module, which aims to jointly model the features from both branches, thereby enhancing the information exchange. (c) represents the ODIM module, which models the temporal attributes of the OD pair features using the Channel Mixer, while the comprehensive relations among OD pairs are learned using the Multi-view Mixer. (d) shows the Channel Mixer, Origin Mixer, Des Mixer.
@@ III Preliminaries
Definition 1. Incomplete OD Matrix $\mathrm{IOD}_{t}^{d}\in\mathbb{R}^{N\times N}$: $\mathrm{IOD}_{t}^{d}(i,j)$ is the number of passengers entering from station $i$ and exiting from station $j$ at the $t^{th}$ time interval of the $d^{th}$ day.
Definition 2. Unfinished OD Matrix $\mathrm{UOD}_{t}^{d}\in\mathbb{R}^{N\times N}$: $\mathrm{UOD}_{t}^{d}(i,j)$ signifies the number of passengers entering station $i$ at the $t^{th}$ time interval and exiting from station $j$ after the $t^{th}$ time interval of the $d^{th}$ day.
Definition 3. Complete OD Matrix $\mathrm{OD}_{t}^{d}\in\mathbb{R}^{N\times N}$: $\mathrm{OD}_{t}^{d}(i,j)$ denotes the number of passengers at the $t^{th}$ time interval entering station $i$ and subsequently exiting from station $j$ at and after the $t^{th}$ time interval of the $d^{th}$ day. $\mathrm{OD}_{t}^{d}=\mathrm{IOD}_{t}^{d}+\mathrm{UOD}_{t}^{d}$.
Definition 4. Unfinished Order Vector $\mathrm{unf}_{t}^{d}\in\mathbb{R}^{N}$: $\mathrm{unf}_{t}^{d}(i)$ represents the number of passengers at the $t^{th}$ time interval entering station $i$ but not exiting from any station at the $t^{th}$ time interval of the $d^{th}$ day. $\mathrm{unf}_{t}^{d}(i)=\sum_{j=1}^{N}\mathrm{UOD}_{t}^{d}(i,j)$.
@@ IV Methodology
$\mathrm{unf}_{t-T+i}^{d}\in\mathbb{R}^{N}$ the unfinished order vector ($i=1,\cdots,n$)
$\mathrm{OD}_{t-T+i}^{d}\in\mathbb{R}^{N\times N}$ the complete OD matrix ($i=1,\cdots,n$)
$\mathrm{UOD}_{t-T+i}^{d}\in\mathbb{R}^{N\times N}$ the unfinished OD matrix ($i=1,\cdots,n$)
$\mathrm{IOD}_{t-T+i}^{d}\in\mathbb{R}^{N\times N}$ the incomplete OD matrix ($i=1,\cdots,n$)
@@ IV-A Architecture Overview
IV-B Dual-branch Structure
As shown in Fig. 2, the OD flow on adjacent days exhibits certain similarities. Therefore, we integrated this similarity into our model by designing a dual-branch structure. The Prev Branch learns historical information, while the Cur Branch learns information at the current time. Since we focus on the similarity between adjacent days, the input to the Prev Branch is the OD matrix from the same time on the previous day. Since yesterday’s order information has been completed, its input is a complete OD matrix and no operation is required. For the Cur branch, the OD matrix is incomplete as some orders have not yet been completed.
If we directly use the incomplete OD matrix, discrepancies between the inputs of the Prev Branch and Cur Branch arise, hindering the model’s ability to effectively perceive complete information for the current time. Therefore, we introduce the OD Matrix Processing module in the Cur branch.
Specifically, we use an Embedding layer to encode the OD matrix based on the OD pair perspective for the inputs of both branches. Subsequently, the ODIM module learns the relations between OD pairs. We then introduce the BTL module to learn the relations between the two branches, supplementing the Cur Branch with historical information. This design captures the similarities between adjacent days and completes the incomplete information for the current time, improving the model’s prediction accuracy and reliability.
This dual-branch structure not only manages the complex relation between OD traffic and time but also enhances the prediction accuracy of current OD traffic by leveraging historical information. Consequently, our model can better handle the large and complex OD matrix in the subway system, utilizing historical data for accurate prediction, thereby demonstrating significant potential and practical value in real-world applications.
IV-C OD Matrix Processing
In this subsection, we introduce the approach for supplementing the OD matrix and estimating the current UOD distribution by leveraging the historical UOD distribution.
IV-C1 Unfinished Orders Processing
@@
Acknowledging the temporal similarity characteristic of metro traffic, where the current OD traffic distribution closely resembles that of the previous time, we estimate the current UOD distribution by referencing historical UOD distributions. Our approach specifically incorporates two temporal scales: long-term and short-term UOD distributions. The calculation is outlined as follows:
${\color[rgb]{0,0,0}\mathrm{US}^{d}_{t-T+i}(j,k)}$ ${\color[rgb]{0,0,0}=}{\color[rgb]{0,0,0}\mathrm{unf}^{d}_{t-T+i}(j)\ast\frac{\mathrm{UOD}^{d-1}_{t-T+i}(j,k)}{\sum_{k=1}^{N}{\mathrm{UOD}^{d-1}_{t-T+i}(j,k)}}}$ (2)
${\color[rgb]{0,0,0}\mathrm{UL}^{d}_{t-T+i}(j,k)}$ ${\color[rgb]{0,0,0}=}{\color[rgb]{0,0,0}\mathrm{unf}^{d}_{t-T+i}(j)\ast\frac{\mathrm{UOD}^{d-7}_{t-T+i}(j,k)}{\sum_{k=1}^{N}{\mathrm{UOD}^{d-7}_{t-T+i}(j,k)}}}$
${\color[rgb]{0,0,0}\widetilde{\mathrm{UOD}}^{d}_{t-T+i}(j,k)}$ ${\color[rgb]{0,0,0}=}{\color[rgb]{0,0,0}\frac{1}{2}(\mathrm{US}^{d}_{t-T+i}(j,k)+\mathrm{UL}^{d}_{t-T+i}(j,k))}$
Here, $i=1,\cdots,T$, $\widetilde{\mathrm{UOD}}^{d}_{t-T+i}$ represents an estimation of unfinished OD matrix at the ${(t-T+i)}^{th}$ time interval of the $d^{th}$ day, the symbol $\ast$ denotes the multiplication of the corresponding elements. $\mathrm{UOD}^{d-1}_{t-T+i}$ represents the short-term UOD distribution, corresponding to the UOD distribution at the same time yesterday, and $\mathrm{UOD}^{d-7}_{t-T+i}$ represents the long-term UOD distribution, aligned with the UOD distribution observed at the corresponding time on the same day of the previous week. This computation aims to generate a more comprehensive and accurate estimate of the UOD distribution at the current time by combining both short-term and long-term UOD distributions.
@@
${\color[rgb]{0,0,0}\widetilde{\mathrm{OD}}^{d}_{t-T+i}(j,k)=\mathrm{IOD}^{d}_{t-T+i}(j,k)+\widetilde{\mathrm{UOD}}^{d}_{t-T+i}(j,k)}$ (3)
where $i=1,\cdots,T$, $\widetilde{\mathrm{OD}}^{d}_{t-T+i}$ is the estimation of complete OD matrix at the current time, $\mathrm{IOD}^{d}_{t-T+i}$ represents the matrix of actual completed orders, and $\widetilde{\mathrm{UOD}}^{d}_{t-T+i}$ denotes the matrix of unfinished orders computed from the historical short-term and long-term UOD distributions. This approach ensures the model receives comprehensive and accurate input information by integrating data from both completed and ongoing orders, thereby enhancing the accuracy of passenger destination predictions.
IV-D Embedding Layer